At the moment of Transit a clock running on universal time ( UT) read 03h 16min 24s . At the previous midnight the sifereal time was 14 h 38 min 54s . Calculate the Right Ascension of the star


1
Expert's answer
2021-11-21T17:26:30-0500

"\alpha=s-t=11^h22^m30^s."